OpenStudy (anonymous):

the digit in the unit place of the product 127^7* 2^24 is

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

sorry the question is 127^79 * 2^24

OpenStudy (anonymous):

no matter

OpenStudy (anonymous):

sorry for the mistake in question

OpenStudy (anonymous):

\(7^0=1,7^1=49,7^3=343,7^4=2401\) so choice of units digit for \(7^{79}\) is \[\{1,9,3\}\]

OpenStudy (anonymous):

as you can see if the exponent is divisible by 4, you will get a 1, but in this case 79 is not divisible by 4, but 80 is, so the units digit of \((127)^{79}\) is 3

OpenStudy (anonymous):

7^1=7 and 7^2=49 is it.....

OpenStudy (anonymous):

yeah that was a typo

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer should be 8

OpenStudy (anonymous):

damn let me start again \[7^1=7,7^2=49,7^3=343,7^4=2401\] conclusion is still the same

OpenStudy (anonymous):

yes, now we have to find units digit for powers of 2 that is easier, choices are \(2,4,8,6\) in that order

OpenStudy (anonymous):

hope that part is clear, \[2^1=2,2^2=4,2^3=8,2^4=16\] and then \(2^5=32\) so they start over

OpenStudy (anonymous):

since 24 is divisible by 4 we know the units digit of \(2^{24}\) is 6

OpenStudy (anonymous):

units digit of \((127)^{79}\) is 3, units digit of \(2^{24}\) is 6 and \(3\times 6=18\) so units digit of the product is 8

OpenStudy (anonymous):

doesn't any one go to parties, shows, bars, dates, friends houses, something on saturday night instead of coming here to do math homework? damn!

OpenStudy (anonymous):

how units digit of (127)^79 is 3

OpenStudy (anonymous):

@experimentX

OpenStudy (kinggeorge):

Let's look at this using modulo 10. \[ 127^7\cdot 2^{24} \pmod{10}\]We can simplify this problem significantly by notice that this is the same as \[ 7^7\cdot 2^{24} \pmod{10}\]Now, using Euler's theorem, and seeing that \(\phi(10)=4\) We can simplify this even further to \[7^3\cdot2^{24}\pmod{10}\]Since 2 isn't coprime to 10, we can't simplify that part. However, as satellite pointed out above, the last digit of \(2^n\) loops from \(2\rightarrow4\rightarrow6\rightarrow8\rightarrow2\), so \(2^{4n}\equiv 2\pmod{10}\) for integers n. Since \(24=4\cdot6\), we can simplify the problem to \[7^3\cdot2\pmod{10}\]This we can actually calculate to be\[3\cdot2\equiv 6\pmod{10}\]So the units digit is 6.

OpenStudy (anonymous):

the answer should be 8

OpenStudy (kinggeorge):

Right. I screwed up with the power of 2. It should be that \(2^{4n}\equiv 6\pmod{10}\) for all \(n>0\). So then you get \[7^3\cdot2^{24}\equiv3\cdot6\equiv8\pmod{10}\]My mistake.

OpenStudy (anonymous):

did nt understand...can u make it clear

OpenStudy (anonymous):

@shameer1 is it not clear from the pattern here \[7^0=1,7^1=49,7^3=343,7^4=2401\] that the units digit of \(127)^{79}\)is 3 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

forget the numbers not in the units place we have \(7^0\to 1, 7^1\to 7,7^2\to 9, 7^3\to 3, 7^4\to 1\) and once we see the one in the units place, the digits will repeat

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that should be more or less clear because it doesn't matter what \(7^4\) is , only that is ends in a 1, so that \(7^5\) will end in a 7 , \(7^6\) will end in a 9, etc

OpenStudy (anonymous):

so once again to sum up, \(7^{79}\) ends in a 3, \(2^{24}\) ends in a 6, and \(3\times 6=18\) so your number ends in an 8